# MERA: A Comprehensive LLM Evaluation in Russian

### v2

@@ 4.2 Scoring
The performance on the tasks is measured with the following metrics (see Tab. 2 for the connection between the tasks and the metrics):
@@
  • The Joint score, or J, is computed following the methodology Logacheva et al. (2022) and is calculated as a combination of three metrics: Style Transfer Accuracy (STA), assessed using a BERT-based classifier; Meaning Preservation Score (SIM), assessed as the cosine similarity of LaBSE sentence embeddings computed between the original text and the model prediction; the naturalness score (FL), assessed using a fluency classifier.
@@
Model Parameters Context length Hugging Face Hub link Paper
Decoder-only
Llama-2-7b
7B
4096
meta-llama/Llama-2-7b-hf Touvron et al. (2023)
Llama-2-13b
13B
4096
meta-llama/Llama-2-13b-hf
Mistral
7B
32768
mistralai/Mistral-7B-v0.1 Jiang et al. (2023)
Yi-6B
6B
4096
01-ai/Yi-6B
ruGPT-3.5
13B
2048
ai-forever/ruGPT-3.5-13B
ruGPT-3-small
125M
2048
ai-forever/rugpt3small_based_on_gpt2 Zmitrovich et al. (2023)
ruGPT-3-medium
355M
2048
ai-forever/rugpt3medium_based_on_gpt2
ruGPT-3-large
760M
2048
ai-forever/rugpt3large_based_on_gpt2
mGPT
1.3B
2048
ai-forever/mGPT Shliazhko et al. (2023)
mGPT-13B
13B
2048
ai-forever/mGPT-13B
Encoder-decoder
FRED-T5-large
820M
512
ai-forever/FRED-T5-large Zmitrovich et al. (2023)
FRED-T5-1.7B
1.7B
512
ai-forever/FRED-T5-1.7B
ruT5-base
 [^2] M
512
ai-forever/ruT5-base
ruT5-large
737M
512
ai-forever/ruT5-large
umT5-Small
300M
512
google/umt5-small Chung et al. (2023)
umT5-Base
580M
512
google/umt5-base
umT5-XL
3.7B
512
google/umt5-xl
umT5-XXL
13B
512
google/umt5-xxl
@@
Calculating overall leaderboard score for aggregation-type benchmarks has faced considerable criticism Rofin et al. (2023). We adopt a methodology aligned with standard scoring systems as demonstrated by Wang et al. (2019); Shavrina et al. (2020b). Our approach involves computing scores for each task based on their metrics by the methodology outlined for the set. Subsequently, the final score is obtained by averaging these task scores, excluding diagnostics tasks from the computation of the final score. In cases where tasks have multiple metrics, these metrics are also averaged. Specifically, for the ruMMLU set, the leaderboard score is averaged across domains internally.
@@ 4.3 Submission
The test sets are not available to anyone except for the organizers, academy, and industry experts supporting the benchmark. The scoring system is automatic and is available on the benchmark platform. The process of submission is the following.
First, users clone MERA benchmark repository [^10] https://github.com/ai-forever/MERA and form submission files using shell script [^11] https://github.com/ai-forever/MERA/blob/main/lm-evaluation-harness/README.md#run-full-benchmark-with-bash-script and the provided customized lm-harness code. Notably, the actual model is not required for submission and evaluation; proprietary models are applicable. Second, they upload the submission files via the platform interface for the automatic assessment. The evaluation result is then displayed in the user’s account and is kept private.
Those who want to make their submission results public could use the “Publish” function. Once activated, the results are forwarded to experts, who validate the submission and verify its reproducibility (all information about the model is provided, the parameters and prompts are intact and were not modified). Once approved, the model’s overall score is shown publicly, while its specific outputs remain private and accessible only to the participating experts.
@@ 5 Baselines
Name
Acc
EM
F1
Acc
Acc
F1 macro
Acc
Acc
Acc
F1 macro
Acc
Acc
F1 macro
Acc
Acc
Llama-2-7b
27.7
1.1
8.1
53.2
34.9
27.2
36.7
12.4
47.5
47.1
50.0
54.5
54.3
50.4
83.9
Llama-2-13b
31.4
1.4
9.8
47.8
32.9
25.8
48.6
15.6
63.7
63.7
49.3
70.3
70.3
50.0
91.1
Mistral
34.4
6.7 12.4
51.8
37.2
34.4
51.6 19.5 73.5 73.2
50.2
81.0 81.1
51.2
95.0
Yi-6B 38.2
5.1
7.9
51.4
33.3
16.7
41.6
18.9
59.0
58.8
50.5
54.1
54.2
49.6
95.1
ruGPT-3.5
25.8
3.6
11.5
50.4
33.1
19.4
0.1
2.5
22.2
20.8
48.8
24.6
22.0
52.3
2.9
ruGPT-3-small
24.4
0.9
6.3
49.8
33.3
16.7
0.1
0.9
25.8
25.3
50.0
25.7
25.4
49.2
0.0
ruGPT-3-medium
24.8
4.3
10.6
49.8
33.3
16.7
0.1
1.2
27.3
27.1
50.0
25.1
24.8
50.0
0.8
ruGPT-3-large
25.1
2.6
9.9
49.8
33.3
16.7
0.1
0.7
21.0
17.8
50.0
23.2
19.1
51.5
0.4
mGPT
25.8
1.4
5.5
49.8
33.3
16.7
0.1
1.2
24.5
19.3
50.0
25.1
22.5
51.9
0.7
mGPT-13B
26.3
2.3
6.2
49.8
33.3
16.7
0.0
1.9
25.0
19.3
50.0
23.2
17.2
48.5
2.3
FRED-T5-large
24.0
0.0
5.2
49.2
35.4
24.8
0.0
0.0
26.5
21.5
49.3
23.2
17.4
49.2
0.0
FRED-T5-1.7B
24.6
0.1
3.1
49.8
33.3
16.7
0.1
0.0
25.0
12.9
49.5
25.5
13.0
50.0
0.0
ruT5-base
25.9
0.0
0.8
50.8
33.6
26.9
0.0
0.0
26.5
18.3
49.3
23.4
15.1
48.1
0.0
ruT5-large
25.4
0.0
1.0
49.8
32.6
29.6
0.0
0.0
26.2
15.8
50.5
25.9
15.9
48.5
0.0
umT5-Small
26.1
0.0
0.3
52.0
28.8
25.5
0.0
0.0
25.5
22.3
50.0
22.5
19.8
48.1
0.0
umT5-Base
25.2
0.0
0.2
46.8
33.6
30.6
0.0
0.0
24.2
14.8
52.6
23.8
14.7
50.0
0.0
umT5-XL
26.1
0.3
1.3
50.6
32.6
18.5
0.0
0.0
23.0
22.3
52.8
26.9
25.5
50.0
0.0
umT5-XXL
24.1
4.1
9.3
49.6
31.5
16.6
0.0
0.0
24.0
16.9
48.8
23.8
14.7
50.4
0.0
Random baseline
24.4
0.1
1.4
48.2
36.1
36.0
0.0
0.0
24.5
24.5
47.2
23.0
22.9
51.9
0.0
Human baseline 99.0 91.0 92.8 98.2 58.7 56.5 99.9 99.8 86.5 87.5 94.2 93.5 93.5 83.5 100.0
Table 4: The results of baseline evaluation on the MERA problem-solving tasks.
@@
Name Acc EM F1
Acc
pass@1
pass@5
pass@10
Acc Grade norm Total score
Llama-2-7b 42.6 0.0 2.1
10.6
0.7
3.4
6.7
45.2
1.4
32.7
Llama-2-13b 50.7 0.0 4.3
9.0
0.8
4.0
7.9
56.3
1.0
36.8
Mistral 39.2 0.0 3.8
9.8
1.2 5.8 11.6 67.6
2.2
40.0
Yi-6B 46.9 0.0 0.8
11.2
0.3
1.5
3.0
48.7
2.3
35.4
ruGPT-3.5 49.2 0.0 3.7
13.2
0.1
0.3
0.6
24.6
2.5
20.8
ruGPT-3-small 36.7 0.0 0.7
8.0
0.0
0.0
0.0
26.3
0.1
19.1
ruGPT-3-medium 43.0 0.0 0.5
10.2
0.0
0.0
0.0
27.1
0.2
20.1
ruGPT-3-large 41.6 0.0 0.7
12.2
0.0
0.0
0.0
24.5
0.0
19.3
mGPT 44.9 0.0 0.4 13.6
0.0
0.0
0.0
24.1
0.0
19.8
mGPT-13B 46.3 0.0 0.6
13.2
0.0
0.0
0.0
23.5
0.2
19.6
FRED-T5-large 47.5 0.0 0.1
8.6
0.0
0.0
0.0
24.8
0.0
19.4
FRED-T5-1.7B 50.8 0.0 0.6
8.8
0.0
0.0
0.0
26.2
0.0
19.1
ruT5-base 48.6 0.0 0.1
10.0
0.0
0.0
0.0
23.7
0.0
19.3
ruT5-large 40.2 0.0 0.0
11.0
0.0
0.0
0.0
24.0
0.0
19.0
umT5-Small 54.6 0.0 0.2
10.8
0.0
0.0
0.0
22.5
0.2
19.8
umT5-Base 52.3 0.0 0.1
10.6
0.0
0.0
0.0
23.1
0.0
19.5
umT5-XL 49.4 0.0 0.1
12.0
0.0
0.0
0.0
25.4
0.1
20.1
umT5-XXL 48.6 0.0 0.3
9.4
0.0
0.0
0.0
29.6
0.4
19.5
Random baseline 50.0 0.0 0.2
9.6
0.0
0.0
0.0
25.8 6.4 20.5
Human baseline 100.0 64.5 71.9 56.0 100.0 100.0 100.0 84.4 70.1 87.2
Table 5: The results of baseline evaluation on the MERA exam-based tasks. We calculate “Total score” from scores of the problem-solving tasks and the exam-based tasks (see Sec. 4.2).

### v3

@@ 4.2 Scoring
The performance on the tasks is measured with the following metrics (see Tab. 2 for the task metrics and the motivation for their choice is given in App. C):
@@
  • The Joint score, or J, is computed following the methodology of Logacheva et al. (2022) and is calculated as a combination of three metrics: Style Transfer Accuracy (STA), assessed using a BERT-based classifier; Meaning Preservation Score (SIM), assessed as the cosine similarity of LaBSE sentence embeddings computed between the original text and the model prediction; the naturalness score (FL), assessed using a fluency classifier.
@@
Model Parameters Context length Hugging Face Hub link Citation
Decoder-only Llama-2-7b 7B 4096 meta-llama/Llama-2-7b-hf Touvron et al. (2023)
Llama-2-13b 13B 4096 meta-llama/Llama-2-13b-hf
Mistral 7B 32768 mistralai/Mistral-7B-v0.1 Jiang et al. (2023)
davinci-002 16384 OpenAI (2024)
Yi-6B 6B 4096 01-ai/Yi-6B Young et al. (2024)
ruGPT-3.5 13B 2048 ai-forever/ruGPT-3.5-13B
ruGPT-3-small 125M 2048 ai-forever/rugpt3small_based_on_gpt2 Zmitrovich et al. (2023)
ruGPT-3-medium 355M 2048 ai-forever/rugpt3medium_based_on_gpt2
ruGPT-3-large 760M 2048 ai-forever/rugpt3large_based_on_gpt2
mGPT 1.3B 2048 ai-forever/mGPT Shliazhko et al. (2024)
mGPT-13B 13B 2048 ai-forever/mGPT-13B
Encoder-decoder FRED-T5-large 820M 512 ai-forever/FRED-T5-large Zmitrovich et al. (2023)
FRED-T5-1.7B 1.7B 512 ai-forever/FRED-T5-1.7B
ruT5-base  [^2] M 512 ai-forever/ruT5-base Zmitrovich et al. (2023)
ruT5-large 737M 512 ai-forever/ruT5-large
umT5-Small 300M 512 google/umt5-small Chung et al. (2023)
umT5-Base 580M 512 google/umt5-base
umT5-XL 3.7B 512 google/umt5-xl
umT5-XXL 13B 512 google/umt5-xxl
@@
Calculating overall leaderboard score for aggregation-type benchmarks has faced considerable criticism Rofin et al. (2023). We adopt a methodology aligned with standard scoring systems as demonstrated by Wang et al. (2019); Shavrina et al. (2020b). For scoring, we first calculate metrics for each task. Then, the final score is computed by averaging these task scores, excluding diagnostics tasks from the computation of the final score. For tasks with multiple metrics, these metrics are also averaged. Specifically, for the ruMMLU set, the leaderboard score is averaged across domains internally.
@@ 4.3 Submission
The test answers are available only for the organizers, and experts supporting the benchmark. The scoring system is automatic and is available on the benchmark platform. The process of submission is the following. First, users clone MERA benchmark repository [^11] https://github.com/ai-forever/MERA and form submission files using shell script [^12] https://github.com/ai-forever/MERA/blob/v1.1.0/lm-evaluation-harness/README.md#run-full-benchmark-with-bash-script and the provided customized lm-eval code. Second, they need to register on the benchmark platform and upload the submission files via the platform interface in their personal account for automatic assessment.
The evaluation result is then displayed in the user’s account and kept private unless they use the “Publish” function and request publication. In this case, it undergoes an expert verification of its reproducibility, which includes checking log files automatically formed by the evaluation script and the provided submission information. Once approved, the model’s score is shown publicly on the leaderboard, while its specific outputs remain private.
@@ 5 Baselines
5.1 Random Baseline
The random baseline is a simple data-agnostic baseline that samples predictions uniformly from the set of target classes in a given task. For most tasks, we randomly choose the result and score the variant. See Sec. E.1 for the details.
5.2 Model Baselines
We evaluated 19 publicly available language models from 10 model families for Russian, including the multilingual ones, varying in size from 125M (ruGPT-3-small) to 13B parameters (Llama-2-13b, and others). See Tab. 3 for the details.
We evaluate models in the same environments and scenarios by the procedure described in Sec. 4.1 and the submission procedure described in Sec. 4.3. See Sec. E.2 for more details.
5.3 Human Baselines
For most tasks, the human evaluation is performed by annotators certified as Russian native speakers via Toloka [^13] https://toloka.ai and ABC [^14] https://elementary.activebc.ru data labeling platforms. See Sec. E.3 for more details.
Human baseline stands for the re-annotation of samples from each task test set through three steps: 1) unpaid training for annotators, 2) paid examination to assess the accuracy of an annotator, and 3) paid main stage to annotate test samples. The annotator is given detailed task instructions, solution criteria, and examples.
The accuracy threshold for the main stage is task-specific and depends on the task difficulty, while the threshold for control tasks on the main equals 50%. The final answer is chosen by majority voting. In the case of the equal answer number, the preference is given to the answer from more skilled annotators. See App. F for other annotation details.
6 Results
The baseline results are summarized in Tab. 4 (problem-solving tasks), Tab. 5 (exam-based tasks), and Tab. 6 (diagnostic tasks) [^15] The version of the code v.1.1.0.. As the evaluation approach is deterministic (see Sec. 4.1), we report results from a single model run.
@@ 6 Results
Name Acc EM F1 Acc Acc F1
macro
EM EM Acc F1
macro
Acc Acc F1
macro
Acc EM
Llama-2-7b 27.7 1.1 8.1 53.2 34.9 27.2 36.7 12.4 47.5 47.1 50.0 54.5 54.3 50.4 83.9
Llama-2-13b 31.4 1.4 9.8 47.8 32.9 25.8 48.6 15.6 63.7 63.7 49.3 70.3 70.3 50.0 91.1
Mistral 34.4 6.7 12.4 51.8 37.2 34.4 51.6 19.5 73.5 73.2 50.2 81.0 81.1 51.2 95.0
davinci-002 35.3 4.4 11.9 50.6 33.1 17.8 47.6 17.6 67.5 67.6 51.9 76.6 76.5 48.1 92.7
Yi-6B 38.2 5.1 7.9 51.4 33.3 16.7 41.6 18.9 59.0 58.8 50.5 54.1 54.2 49.6 95.1
ruGPT-3.5 25.8 3.6 11.5 50.4 33.1 19.4 0.1 2.5 22.2 20.8 48.8 24.6 22.0 52.3 2.9
ruGPT-3-small 24.4 0.9 6.3 49.8 33.3 16.7 0.1 0.9 25.8 25.3 50.0 25.7 25.4 49.2 0.0
ruGPT-3-medium 24.8 4.3 10.6 49.8 33.3 16.7 0.1 1.2 27.3 27.1 50.0 25.1 24.8 50.0 0.8
ruGPT-3-large 25.1 2.6 9.9 49.8 33.3 16.7 0.1 0.7 21.0 17.8 50.0 23.2 19.1 51.5 0.4
mGPT 25.8 1.4 5.5 49.8 33.3 16.7 0.1 1.2 24.5 19.3 50.0 25.1 22.5 51.9 0.7
mGPT-13B 26.3 2.3 6.2 49.8 33.3 16.7 0.0 1.9 25.0 19.3 50.0 23.2 17.2 48.5 2.3
FRED-T5-large 24.0 0.0 5.2 49.2 35.4 24.8 0.0 0.0 26.5 21.5 49.3 23.2 17.4 49.2 0.0
FRED-T5-1.7B 24.6 0.1 3.1 49.8 33.3 16.7 0.1 0.0 25.0 12.9 49.5 25.5 13.0 50.0 0.0
ruT5-base 25.9 0.0 0.8 50.8 33.6 26.9 0.0 0.0 26.5 18.3 49.3 23.4 15.1 48.1 0.0
ruT5-large 25.4 0.0 1.0 49.8 32.6 29.6 0.0 0.0 26.2 15.8 50.5 25.9 15.9 48.5 0.0
umT5-Small 26.1 0.0 0.3 52.0 28.8 25.5 0.0 0.0 25.5 22.3 50.0 22.5 19.8 48.1 0.0
umT5-Base 25.2 0.0 0.2 46.8 33.6 30.6 0.0 0.0 24.2 14.8 52.6 23.8 14.7 50.0 0.0
umT5-XL 26.1 0.3 1.3 50.6 32.6 18.5 0.0 0.0 23.0 22.3 52.8 26.9 25.5 50.0 0.0
umT5-XXL 24.1 4.1 9.3 49.6 31.5 16.6 0.0 0.0 24.0 16.9 48.8 23.8 14.7 50.4 0.0
Random baseline 24.4 0.1 1.4 48.2 36.1 36.0 0.0 0.0 24.5 24.5 47.2 23.0 22.9 51.9 0.0
Human baseline 99.5 91.0 92.8 98.2 61.0 58.3 99.9 100.0 86.5 87.5 94.2 83.7 83.8 83.8 100.0
Table 4: The results of baseline evaluation on the MERA problem-solving tasks. Best model scores are underlined.
@@
Name Acc EM F1 Acc pass@1 pass@5 pass@10 Acc Grade norm Total score
Llama-2-7b 42.6 0.0 2.1 10.6 0.7 3.4 6.7 45.2 1.4 32.7
Llama-2-13b 50.7 0.0 4.3 9.0 0.8 4.0 7.9 56.3 1.0 36.8
Mistral 39.2 0.0 3.8 9.8 1.2 5.8 11.6 67.6 2.2 40.0
davinci-002 52.1 0.0 1.8 12.4 0.5 2.3 3.7 61.3 1.6 38.3
Yi-6B 46.9 0.0 0.8 11.2 0.3 1.5 3.0 48.7 2.3 35.4
ruGPT-3.5 49.2 0.0 3.7 13.2 0.1 0.3 0.6 24.6 2.5 20.8
ruGPT-3-small 36.7 0.0 0.7 8.0 0.0 0.0 0.0 26.3 0.1 19.1
ruGPT-3-medium 43.0 0.0 0.5 10.2 0.0 0.0 0.0 27.1 0.2 20.1
ruGPT-3-large 41.6 0.0 0.7 12.2 0.0 0.0 0.0 24.5 0.0 19.3
mGPT 44.9 0.0 0.4 13.6 0.0 0.0 0.0 24.1 0.0 19.8
mGPT-13B 46.3 0.0 0.6 13.2 0.0 0.0 0.0 23.5 0.2 19.6
FRED-T5-large 47.5 0.0 0.1 8.6 0.0 0.0 0.0 24.8 0.0 19.4
FRED-T5-1.7B 50.8 0.0 0.6 8.8 0.0 0.0 0.0 26.2 0.0 19.1
ruT5-base 48.6 0.0 0.1 10.0 0.0 0.0 0.0 23.7 0.0 19.3
ruT5-large 40.2 0.0 0.0 11.0 0.0 0.0 0.0 24.0 0.0 19.0
umT5-Small 54.6 0.0 0.2 10.8 0.0 0.0 0.0 22.5 0.2 19.8
umT5-Base 52.3 0.0 0.1 10.6 0.0 0.0 0.0 23.1 0.0 19.5
umT5-XL 49.4 0.0 0.1 12.0 0.0 0.0 0.0 25.4 0.1 20.1
umT5-XXL 48.6 0.0 0.3 9.4 0.0 0.0 0.0 29.6 0.4 19.5
Random baseline 50.0 0.0 0.2 9.6 0.0 0.0 0.0 25.8 6.4 20.5
Human baseline 100.0 64.5 71.9 70.4 100.0 100.0 100.0 89.0 70.1 88.7
Table 5: The results of baseline evaluation on the MERA exam-based tasks. “Total score” is computed based on scores of the problem-solving tasks and the exam-based tasks (see Sec. 4.2). Best model scores are underlined.